OpenStudy (anonymous):

Write the equation of the line that has an x-intercept of -3 and passes through the point (-3, 7)

OpenStudy (anonymous):

x=-3

OpenStudy (anonymous):

unless you meant the y-intercept. as of how it is written now, it is a vertical line

OpenStudy (anonymous):

how did you find that?

OpenStudy (anonymous):

x-intercept means it crosses the point (-3,0) the other point you have is (-3,7) which is 2 values for y when x=-3 the slope of the line is undefined(infinity) has to be vertical line.

OpenStudy (anonymous):

slope = rise/run = y1-y2 0-(-7) +7 ------ = ------- = ------- x1-x2 -3-(-3) 0